# Reinforcement Twinning: from digital twins to model-based reinforcement learning

### 31 Jan 2024 v2

@@ 6.3 Trajectory control of flapping wings micro air vehicles
The training detailed in Section 6.1 converges to one optimal policy that the drone applies to control its flight in the real environment. Figure 18 characterizes one such flight depicting the state variables (17(a) and 17(b)) and the control actions (17(c)). The figures evidence that the drone managed to closely reach its target ($\bm{x}_{f}=5$ m and $\bm{\dot{x}}_{f}=0$, as fast as possible) thanks to the control of the stroke plane angle $\beta_{w}$ and the flapping amplitude $A_{\phi}$. The control of the latter is mainly formed of three phases in a kind of bang-bang fashion. First, the drone flaps its wings at their maximum angle of attack ($A_{\phi}=88$ deg) to reach the target position as fast as possible. Then, the drone picks a flapping angle close to the minimal one ($55$ deg) to decelerate and stabilize at the target height with minimal overshoot. Finally, the wings flap at a slightly bigger angle so that they generate enough lift to balance the body weight. The lift is also slightly modulated by the other control variable ($\beta_{w}$) which mainly drives the horizontal motion of the drone. At the start of the episode, the drone pitches the stroke plane up to the maximum angle ($\beta_{w}=-30$ deg) and then stabilizes it around an angle that allows it to stay close to the desired position, $x_{f}$, despite the headwind. The influence of the gust is visible on $\beta_{w}$ even though the wind effect is still very limited, due to the selected mass and surface of the body.
@@
Figure 19 analyses the policies found by the model-free (Figures 18(a) - 18(d)) and model-based (Figures 18(e) - 18(h)) loops. To this end, we fix two of the four inputs to zero and span on the other two dimensions, recording the corresponding action. The two policies are remarkably similar, converging to essentially the same parametrization with their differences lying in the concavity of the hyperbolic tangent, as dictated by the biases. This affects the localization of the switching between the two bounds of the actuation space, possibly explaining the differences in their associated costs presented in Table 2. In both cases, for what concerns $\hat{A}_{\phi}$, the demarcation zone between the two bounds lies on the diagonal that goes from $(\hat{\dot{e}}_{x},\hat{e}_{x})=(-1.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x},)=(0.25,1.0)$, see Figure 18(a). The same trend is observed in Figure 18(b), which analyses the dependencies from $e_{z}$ and its time derivative. In this case, however, the upper point of the diagonal lies around $(\hat{\dot{e}}_{z},\hat{e}_{z})=(0.0,1.0)$, thus increasing $\hat{A}_{\phi}$ as the rate of change of $e_{z}$ goes negative, and decreasing it otherwise. Interestingly, the model-based policy here shows a broader decision zone, as depicted in Figure 9(f). Now, analysing the actuation on $\beta_{w}$, the threshold appears to be positioned in the diagonals $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.75,1.0)$. The sensitivity to the $z-$errors instead show it to lie in $(\hat{\dot{e}}_{z},\hat{e}_{z})=(-1.0,0.0)$ and $(\hat{\dot{e}}_{z},\hat{e}_{z})=(1,0.5)$. Therefore, it seems that the time variation of the rate of change is the most important input on the $x-$dimension, whereas in the $z-$coordinate, it is the scalar $\hat{e}_{z}$. Arguably this may be caused by the headwind, oriented as opposite of the $x-$versor. Overall, there appears to be two main mechanisms at play: (1) the errors and their rate of change seems to have the same importance when deciding how to control on $\hat{A}_{\phi}$ being slightly more sensitive to the scalar instantaneous errors; and (2) when deciding on how to actuate $\beta_{w}$, the controller balances the scalar instantaneous error in one dimension and its rate of change in the other. In fact, when $\dot{e}_{x}$ increases, it actuates on $\beta_{w}$ to get back to nominal conditions, and the same is observed on analysing the reactivity to the instantaneous error $e_{z}$ and $e_{x}$.
@@
Figure 19: Comparison of model-based and model-free policies
We conclude the FWMAV analysis presenting the robustness study of the model-free and model-based policies, shown in Figure 20. One-hundred episodes were run on the real environment for each policy and their costs are compared with violins plots (Figure 19(a)). Both sets of weights result in a small cost proving their capability to handle the random gust. The model-free policy manages to reach on average a lower cost and shows a slightly different distribution than the model-based policy, having two peaks at $\mathcal{L}_{a}=905$ and $\mathcal{L}_{a}=920$, whereas the model-based shows a main peak in the middle, at $L_{a}=910$. However, their differences are in the order $\mathcal{O}(10^{1})$, reinforcing the observation that the two have converged to a very similar policy. Lastly, The same procedure is then applied on the assimilation side, studying its robustness to varying operative conditions, as shown in Figure 19(b). It appears that the identified closure function is robust to different operative conditions, registering a mean cost of $\mathcal{L}_{p}\approx\mathcal{O}(10^{-3})$.
@@
6.4 Thermal management of cryogenic storage
@@
The control performances of the optimal parametrization, $\bm{w}_{a}^{*}$, extracted from the learning curve presented in Figure 9(d) are studied next. The current scenario involves an initial over-pressurized tank operating 0.4 bar above the reference $\tilde{p}$. Figure 22 shows the time evolution of the real pressure $\check{p}$ and propellant mass $\check{m}_{l}$ over time, while the vented mass flow rate $\check{\dot{m}}_{c}$. Figures 21(b) and 21(c) show the control action over time and the corresponding pressure, respectively. Since the tank is initially in an over-pressurized state sufficiently far from $\tilde{p}$, the valve was regulated to open to at its widest ($\dot{m}_{c}=0.63$ g/s) allowing the pressure to rapidly reduce until the target 5 bar are reached. This pressure drop is accompanied by a consumption of 87 kg over two roughly 45 hours. Once the target is reached, the active TVS regulates the pressure around this value. To do so, $\check{\dot{m}}_{c}$ fluctuates between -50% and 25% around its midpoint due to changes in the heat due to the disturbance, although such variations are not seen in the pressure signal.
@@
From these findings, we conclude that the control policy is able to completely control the tank pressure around its desired target value. However, the response time of the TVS is limited by the technical constraints assigned to the circulating pump, and Joule-Thomson valve. In addition, depending on the value assigned to the $K_{1}$ constant in 47, one could retrieve either faster or more fuel-efficient policies.
@@
We inspect the two found policies in Figure 23. The two loops, model-based and model-free, converge to essentially the same parametrization. They show almost identical response to the possible input space, with very minor discrepancies in the inclination of the hyperbolic tangent. Lastly, the robustness of the assimilation and control loops concludes this overview. The procedure consists in evaluating the cost of the optimal weights $\bm{w}_{p}$ and $\bm{w}_{a}$ over observations unseen during the training phase. To this end, we carried out one-hundred simulations on random initial conditions and disturbances. For each of them, we computed the assimilation cost $\mathcal{L}_{p}$, as well as the control cost $\mathcal{L}_{a}$ using the model-based and model-free policies. This investigation, summarized in the violin plots presented in Figures 23(a) and 23(b), proves that the weights obtained at the end of the learning phase are valid and robust even in unseen scenarios. Moreover, in contrast to what was observed in the wind turbine test case, the model-based and model-free weights are nearly identical, thus their performances are also identical as supported by Table 2 which shows a difference between the two control performances in the order of $\mathcal{O}(10^{-2})$. This is likely a consequence of both approaches converging in the same region of the search space.
@@ 7 Conclusions and outlooks
We presented a novel framework to train physics-based digital twins from real-time data while at the same time training a control agent through a combination of model-free and model-based methods. We called this framework Reinforcement Twinning (RT), and we positioned it at the intersection between data assimilation, system identification and reinforcement learning. The digital twins are formulated as physics-based models relying on closure laws expressed as parametric functions. The parameters defining these closures are identified from real-time data using and adjoint-based data assimilation strategy.
We presented a novel framework to train physics-based digital twins from real-time data while at the same time training a control agent through a combination of model-free and model-based methods. We called this framework Reinforcement Twinning (RT) and positioned it at the intersection between data assimilation, system identification and reinforcement learning. The digital twins are formulated as physics-based models relying on closure laws expressed as parametric functions. The parameters defining these closures are identified from real-time data using an adjoint-based data assimilation strategy.
@@
We tested this novel framework on three widely different engineering systems: a wind turbine in a time-varying wind speed, a flapping-wing micro air vehicle travelling against gusts, and the thermal management of a cryogenic tank subject to time-varying heat loads. In these problems, the physics-based model of the digital twin is formulated as ODEs ensuring the compliance of predictions with fundamental principles (conservation of angular momentum in a wind turbine, Newton’s law in the drone’s trajectory and conservation of mass/energy within the cryogenic tank) and the closure law describe unmodelled physics (aerodynamic coefficients and heat transfer correlations).

### 25 Feb 2024 v3

@@ 6.3 Performances on a FWMA Trajectory Control
Figure 18 characterizes the identified optimal policy, depicting state variables (17(a) and 17(b)) and control actions (17(c)). The figures evidence that the drone managed to closely reach its target ($\bm{x}_{f}=5$ m and $\bm{\dot{x}}_{f}=0$, as fast as possible) thanks to the control of the stroke plane angle $\beta_{w}$ and the flapping amplitude $A_{\phi}$. The control of the latter is mainly formed of three phases in a kind of bang-bang fashion. First, the controller sets the drone’s wings at their maximum angle of attack ($A_{\phi}=88$ deg) to reach the target position as fast as possible. Then, it decelerates and stabilizes the drone picking a flapping angle close to the minimal one ($55$ deg), with a minimal overshoot. Finally, the wings’ angle of attack is increased again, to balance the body weight with the aerodynamic lift. This force is also slightly modulated by the other control variable ($\beta_{w}$) which mainly drives the horizontal motion of the drone. In the same way, the controller pitches the stroke plane up to the maximum angle ($\beta_{w}=-30$ deg) and then stabilizes it around an angle that allows it to stay close to the desired position, $x_{f}$, despite the headwind. The influence of the gust is visible on $\beta_{w}$ even though the wind effect is still very limited, due to the selected mass and surface of the body.
@@
Figure 19 analyses the policies found by the model-free (Figures 18(a) - 18(d)) and model-based (Figures 18(e) - 18(h)) loops. To this end, we fix two of the four inputs to zero and span on the other two dimensions, recording the corresponding action. The two policies are remarkably similar, converging to essentially the same parametrization. The main difference between the two lies in the concavity of the hyperbolic tangent, as dictated by the biases. This affects the localization of the switching between the two bounds of the actuation space, possibly explaining the differences in their associated costs presented in Table 2. In both cases, for what concerns $\hat{A}_{\phi}$, the border between the two bounds lies on the diagonal that goes from $(\hat{\dot{e}}_{x},\hat{e}_{x})=(-1.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x},)=(0.25,1.0)$, see Figure 18(a). The same trend is observed in Figure 18(b), which analyses the dependencies from $e_{z}$ and its time derivative. In this case, however, the upper point of the diagonal lies around $(\hat{\dot{e}}_{z},\hat{e}_{z})=(0.0,1.0)$, thus increasing $\hat{A}_{\phi}$ as the rate of change of $e_{z}$ goes negative, and decreasing it otherwise. Interestingly, the model-based policy here shows a broader decision zone, as depicted in Figure 9(f). Now, analysing the actuation on $\beta_{w}$, the threshold appears to be positioned in the diagonals $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.0,-1.0)$ to $(\hat{\dot{e}}_{x},\hat{e}_{x})=(0.75,1.0)$. The sensitivity to the $z-$errors instead show it to lie in $(\hat{\dot{e}}_{z},\hat{e}_{z})=(-1.0,0.0)$ and $(\hat{\dot{e}}_{z},\hat{e}_{z})=(1,0.5)$. Therefore, it seems that the time variation of the rate of change is the most important input on the $x-$dimension, whereas in the $z-$coordinate, it is the scalar $\hat{e}_{z}$. Arguably this may be caused by the headwind, oriented as opposite of the $x-$axis. Overall, there appears to be two main mechanisms at play: (1) the errors and their rate of change seems to have the same importance when deciding how to control on $\hat{A}_{\phi}$ being slightly more sensitive to the scalar instantaneous errors; and (2) when deciding on how to actuate $\beta_{w}$, the controller balances the scalar instantaneous error in one dimension and its rate of change in the other. In fact, when $\dot{e}_{x}$ increases, it actuates on $\beta_{w}$ to get back to nominal conditions, and the same is observed on analysing the reactivity to the instantaneous error $e_{z}$ and $e_{x}$.
@@
Figure 19: Comparison of model-based and model-free policies.
We conclude the analysis of the RT algorithm performances on the FWMAV presenting the robustness study of the model-free and model-based policies, shown in Figure 20. One-hundred episodes were run on the real environment for each policy and their costs are compared with violin plots (Figure 19(a)). Both sets of weights result in a small cost proving their capability to handle the random gust. The model-free policy manages to reach on average a lower cost and shows a slightly different distribution than the model-based policy, having two peaks at $\mathcal{L}_{a}=905$ and $\mathcal{L}_{a}=920$, whereas the model-based shows a main peak in the middle, at $\mathcal{L}_{a}=910$. However, their differences are in the order $\mathcal{O}(10^{1})$, reinforcing the observation that the two have converged to a very similar policy. Lastly, the same procedure is applied on the assimilation side, studying its robustness to varying operative conditions, as shown in Figure 19(b). It appears that the identified closure function is robust to different operative conditions, registering a mean cost of $\mathcal{L}_{p}\approx\mathcal{O}(10^{-3})$.
@@
6.4 Performances in Cryogenic Storage’s Thermal Management
@@
The control performances of the optimal parametrization, $\bm{w}_{a}^{*}$, extracted from the learning curve presented in Figure 9(d) are studied next. The current scenario involves an initial over-pressurized tank operating 0.4 bar above the reference $\tilde{p}$. Figure 22 shows the time evolution of the real pressure $\check{p}$ and propellant mass $\check{m}_{l}$ over time, while the vented mass flow rate $\check{\dot{m}}_{c}$. Figures 21(b) and 21(c) show the control action over time and the corresponding pressure, respectively. Since the tank is initially in an over-pressurized state sufficiently far from $\tilde{p}$, the valve was regulated to open to its widest ($\dot{m}_{c}=0.63$ g/s), allowing the pressure to rapidly reduce until the target 5 bar are reached. This pressure drop is accompanied by a consumption of 87 kg over two roughly 45 hours. Once the target is reached, the active TVS regulates the pressure around this value. To do so, $\check{\dot{m}}_{c}$ fluctuates between -50% and 25% around its midpoint due to changes in the heat due to the disturbance, although such variations are not seen in the pressure signal.
@@
From these findings, we conclude that the control policy is able to completely control the tank pressure around its desired target value. However, the response time of the TVS is limited by the technical constraints assigned to the circulating pump and Joule-Thomson valve. In addition, depending on the value assigned to the $K_{1}$ constant in 47, one could retrieve either faster or more fuel-efficient policies.
@@
We inspect the two found policies in Figure 23. The two loops, model-based and model-free, converge to essentially the same parametrization. They show almost identical responses to the possible input space, with very minor discrepancies in the inclination of the hyperbolic tangent. Lastly, the robustness of the assimilation and control loops concludes this overview. The procedure consists in evaluating the cost of the optimal weights $\bm{w}_{p}$ and $\bm{w}_{a}$ over observations unseen during the training phase. To this end, we carried out one hundred simulations on random initial conditions and disturbances. For each of them, we computed the assimilation cost $\mathcal{L}_{p}$, as well as the control cost $\mathcal{L}_{a}$ using the model-based and model-free policies. This investigation, summarized in the violin plots presented in Figures 23(a) and 23(b), proves that the weights obtained at the end of the learning phase are valid and robust even in unseen scenarios. Moreover, in contrast to what was observed in the wind turbine test case, the model-based and model-free weights are nearly identical, thus their performances are also identical as supported by Table 2, which shows a difference between the two control performances in the order of $\mathcal{O}(10^{-2})$. This is likely a consequence of both approaches converging in the same region of the search space.
@@ 7 Conclusions and Outlooks
We presented a novel framework to train physics-based digital twins from real-time data while at the same time training a control agent through a combination of model-free and model-based methods. We called this framework Reinforcement Twinning (RT) and positioned it at the intersection between data assimilation, system identification, and reinforcement learning. The digital twins are formulated as physics-based models relying on closure laws expressed as parametric functions. The parameters defining these closures are identified from real-time data using an adjoint-based data assimilation strategy.
@@
We tested this novel framework on three widely different engineering systems: a wind turbine in a time-varying wind speed, a flapping-wing micro air vehicle traveling against gusts, and the thermal management of a cryogenic tank subject to time-varying heat loads. In these problems, the physics-based model of the digital twin is formulated as ODEs ensuring the compliance of predictions with fundamental principles (conservation of angular momentum in a wind turbine, Newton’s law in the drone’s trajectory, and conservation of mass/energy within the cryogenic tank) and the closure law describe unmodelled physics (aerodynamic coefficients and heat transfer correlations).
@@
Acknowledgements
L. Schena and R. Poletti are supported by Fonds Wetenschappelijk Onderzoek (FWO), while P. Marques and J. van Den Berghe are supported by the FRIA grant from the ‘Fonds de la Recherche Scientifique (F.R.S. -FNRS)’. S.Ahizi and M.A. Mendez were supported by the Flemish Agentschap Innoveren & Ondernemen in the framework of the CSBO project “Clean Hydrogen Propulsion for Ships (CHyPS)”.